# L-WISE: Boosting human image category learning through model-based image selection and enhancement

### v1

@@ Abstract
The currently leading artificial neural network (ANN) models of the visual ventral stream – which are derived from a combination of performance optimization and robustification methods -– have demonstrated a remarkable degree of behavioral alignment with humans on visual categorization tasks. Extending upon previous work, we show that not only can these models guide image perturbations that change the induced human category percepts, but they also can enhance human ability to accurately report the original ground truth. Furthermore, we find that the same models can also be used out-of-the-box to predict the proportion of correct human responses to individual images, providing a simple, human-aligned estimator of the relative difficulty of each image. Motivated by these observations, we propose to augment visual learning in humans in a way that improves human categorization accuracy at test time. Our learning augmentation approach consists of (i) selecting images based on their model-estimated recognition difficulty, and (ii) using image perturbations that aid recognition for novice learners. We find that combining these model-based strategies gives rise to test-time categorization accuracy gains of 33-72% relative to control subjects without these interventions, despite using the same number of training feedback trials. Surprisingly, beyond the accuracy gain, the training time for the augmented learning group was also shorter by 20-23%. We demonstrate the efficacy of our approach in a fine-grained categorization task with natural images, as well as tasks in two clinically relevant image domains – histology and dermoscopy – where visual learning is notoriously challenging. To the best of our knowledge, this is the first application of ANNs to increase visual learning performance in humans by enhancing category-specific features.
@@ 1 Introduction
Figure 1: Robustified ANNs can be used out-of-the-box as image-recognition difficulty estimators and ground-truth percept enhancers. We consider a 16-way basic animal classification task: A1 Correspondence between human categorization accuracy and the model-computed ground-truth logit activation value. The curve denotes a logistic regression model predicting the probability of a correct response based on the logit value (AUC=0.71 under 10-fold cross validation, $p<0.001$). A2 Example images with varying ground truth logit values (predicted difficulty) B1 Perturbing images by ground truth logit maximization increases recognition accuracy progressively with the perturbation $\ell_{2}$-norm pixel budget $\epsilon$. Other off-the-shelf image enhancement methods do not increase categorization accuracy, despite inducing larger perturbations of $\epsilon=43$, $\epsilon=106$, and $\epsilon=26$ on average from left to right. B2 Example images: Unmodified (left), enhanced by ground truth logit maximization with pixel budget ${\epsilon}=10$ and ${\epsilon}=20$, and by off-the-shelf methods (right). All vertical error bars are 95% CI by bootstrap. Horizontal error bars in A1 denote SD over images within each of the logit value bins.
@@
Figure 2: Robustified ANNs can be used to boost novel-object recognition learning in humans. A novice human learner performs an image categorization task, which consists of a training phase (B) and a test phase (C). Images for both phases are randomly drawn from a labeled image dataset of fine-grained unfamiliar categories (A). Feedback (correct/incorrect, with indication of the correct category) is delivered during the training phase only. Our proposed “Logit-Weighted Image Selection and Enhancement” (L-WISE) approach uses an ANN model (D) to augment the visual learning curriculum by using the difficulty score to sample images based on a predefined profile of maximal difficulty per trial (E), and by enhancing images for easier recognition with an enhancement magnitude of a predefined trial-sequence profile (F).
@@
Over the last decade, specific artificial neural network (ANN) models have been shown to be the best image-computable emulators of neural processing along the human and monkey ventral visual stream and its support of a range of human visual tasks, as measured by human behavior. Iterative efforts have developed even better models in this same vein that are more and more accurate emulators. Indeed, among the most contemporary of such models – so called “robustified” deep ANN models (Ma̧dry et al., 2018) – have been shown to allow the design of images to predictably control both ventral stream neural activity (Guo et al., 2022) and human object categorization reports (Gaziv et al., 2024; Croce & Hein, 2020).
Learning to recognize new, unfamiliar categories in images is a task that the human visual system normally excels at. Importantly, despite being a basic task for humans, visual learning of new categories often carries significant practical relevance: for example, medical students and residents devote numerous hours to mastering the diagnosis of various diseases in image modalities such as histology (microscopic images of cells and tissues) and dermoscopy (skin lesions). They achieve this by practicing category recognition on a variety of imaged examples representing different cases. Studies of perceptual learning with simple tasks, such as line orientation discrimination, effectively show a curriculum effect whereby providing easy trials to a novice learner before gradually increasing the difficulty promotes faster perceptual learning (Lu & Dosher, 2022). Motivated by these findings, here, we asked whether the human-behavioral and neural alignment that is attributed to these models makes them useful in helping humans learn to recognize new, unknown categories. Enhancing learning with a model, therefore, serves as both another scientific test of the model’s validity and a beneficial application for human education.
@@
To test the viability of model-guided boosting of visual category learning in humans, we first established two key empirical observations, summarized in Fig. 1: (i) We found that the human error rate in an object categorization task is highly predicted by the ground truth logit activation of a robustified ANN, rendering it a valid image-recognition difficulty score for humans. (ii) We found that this relationship also holds in reverse – pixel-level perturbations can be guided using the model in a way that increases the ground truth logit activation, generating a “perturbed” version of the image that is “easier” to recognize as the ground truth label suggests. Notably, other off-the-shelf image enhancement methods do not result in significant increases in categorization accuracy by humans. We thus propose a method that combines model-based image difficulty prediction and enhancement to generate optimized curricula for novice humans learning challenging image classification tasks.
@@
Our proposed method, “Logit-Weighted Image Selection and Enhancement” (L-WISE), is illustrated in Fig. 2. In the general setting, a novice human learner completes a categorization learning task on an online platform. A labeled image dataset, with unfamiliar category labels, is used to teach the categories by showing examples and providing per-trial feedback after each image category judgment by the learner. Upon completion of the training phase, test accuracy is measured on held-out images in similar trials without feedback (Fig. 2ABC).
L-WISE intervenes on this naive visual learning baseline using a robustified ANN model in two ways: (i) it uses the difficulty score to sample images based on a predefined profile of maximal difficulty per trial (Fig. 2DE); (ii) it “perturbs” training images to be easier for recognition with an enhancement magnitude parameter similarly governed by a predefined trial-sequence profile (Fig. 2D-F).
@@
Surprisingly, despite the human visual system being well-adapted for rapidly learning new visual categories, we find that L-WISE gives rise to substantial gains of 33-72% in test-time accuracy margins above chance relative to control subjects without these interventions. In addition to improved accuracy, the duration of the training for the augmented learning group was significantly reduced. We further demonstrate the practical utility of our approach in boosting visual category learning across varied image domains and category spaces. In particular, we considered three main categorization tasks: moth species in natural images, dermoscopy images, and histology images.
Notably, our approach is inspired by recent findings demonstrating model-guided perturbations that accurately modulate category percepts away from the ground truth label (Gaziv et al., 2024). In this work, we conversely seek to amplify the ground truth percept, i.e., to facilitate the correct categorization of a given image.
Our contributions are several-fold:
@@
 $\bullet$ We establish a new state-of-the-art in predicting image recognition difficulty for humans, using a simple approach employing robustified ANN logit activations.
@@
 $\bullet$ We show that the leading models of the ventral visual stream can guide image perturbations that enhance human ability to accurately report the original ground truth label.
@@
 $\bullet$ We propose a novel model-based visual learning augmentation approach for humans that substantially increases test-time categorization accuracy at a reduced training time. To the best of our knowledge, this is the first application of image enhancement to augment human visual learning.
@@
 $\bullet$ We demonstrate the general applicability of our proposed approach on categorization tasks across a range of image domains – including clinically relevant ones.
@@ 2 Related Work
We develop two important capabilities that form the foundation of our approach to assisting learners: (1) state-of-the-art predictions of the classification difficulty of images for humans, and (2) image perturbations that increase human categorization accuracy. Many works have ranked the difficulty of images to design curricula for ANNs (Wang et al., 2021b). Leading approaches include the c-score learning speed proxy (Jiang et al., 2021) and the prediction depth (Baldock et al., 2021) calculated for each image. Mayo et al. (2023) applied these techniques to predict the difficulty of natural images for humans, defined as either the minimum viewing time required to classify a given image correctly, or (as in our work) the proportion of humans who correctly classify it. Here, we show that the logit score associated with the ground truth class from a robustified ANN model more accurately predicts human image difficulty than prior approaches.
@@
Enhancing image quality has been the focus of many previous studies (Qi et al., 2021), ranging from correction of factors such as lighting and contrast (e.g., Zuiderveld (1994); Jobson et al. (1997)) to ANN models that “upsample” images to higher resolution (Anwar et al., 2020). However, very little research has focused on enhancing images to more strongly represent a specific category. Prior works in this vein focused on making images easier for ANN models to classify correctly (Kim et al., 2023; Tussupov et al., 2023) or less vulnerable to subsequent adversarial attacks (Salman et al., 2021; Frosio & Kautz, 2023). Such perturbations, however, do not strongly affect human perception due to misalignment between human and machine perception (Gaziv et al., 2024).
@@
Other studies focused on model-human alignment. Brain-Score directly benchmarks ANN models with respect to neural representation and downstream behavior (Schrimpf et al., 2018); “Harmonization” methods directly drive alignment by an auxiliary objective on ANN-predicted feature importance maps and crowd-sourced ones (Fel et al., 2022). Other works introduce architecture components to account for additional aspects of human vision, such as the dorsal-stream “where” pathway in the brain (Choi et al., 2023).
@@
A key property that enables ANNs to generate human-interpretable image perturbations is that of perceptually aligned gradients, which is closely related to adversarial robustness and can be induced through adversarial training Ganz et al. (2023); Gaziv et al. (2024). Here, we apply adversarially-trained ANNs to enhance images such that they are more strongly associated with their ground truth label by the guiding model and by humans. To the best of our knowledge, we are the first to demonstrate improved human performance on image classification tasks through category-specific image enhancement.
@@
Our primary goal is to apply difficulty prediction and image enhancement to assist human learning. The emerging field of machine teaching (Zhu, 2015) employs machine learning to find or generate optimal “teaching sets” that can be used to train other models or humans. While many such approaches have been successfully applied to training machine learning models (e.g., Liu et al. (2017); Qiu et al. (2023)), few studies have successfully enhanced image category learning in humans and most of these focus on teaching set selection. Singla et al. (2014) propose STRICT, which optimizes the expected decrease in learner error based on how the selected images and their labels constrain a linear hypothesis class in a feature space. Johns et al. (2015) extend a similar approach to select images in an online fashion by modeling the learner’s progress. MaxGrad (Wang et al., 2021a) uses bi-level optimization to iteratively refine a teaching set by modeling learners as optimal empirical risk minimizers. Most similar to our work are approaches like EXPLAIN (Mac Aodha et al., 2018), which uses ANN class activation maps (CAMs) to highlight relevant image regions while providing feedback to the learner. EXPLAIN also selects a curriculum of images based on (a) a multi-class adaptation of STRICT, (b) representativeness (mean feature-space distance to other images of the same class), and (c) the estimated difficulty (entropy) of the CAM explanations. Chang et al. (2023) use bounding boxes to highlight image regions attended to by experts and not novices, allowing humans to more accurately match bird or flower images to one species among five shown in a gallery.
@@
Our approach departs from previous studies in several ways. We make explicit estimates of image difficulty with unprecedented accuracy to select easier images for early-stage learners. We are unique in employing category-specific image enhancement, which is a novel technique in itself, to improve the teaching efficacy of a given set of images. While Mac Aodha et al. (2018) and Chang et al. (2023) help learners by explicitly highlighting where learners should attend to in the image, we take a distinct and complementary approach by implicitly highlighting what learners must attend to in order to classify images correctly.
@@ 3 Overview of approach and experiments
Our approach to improve visual learning in humans is based on two key observations in the leading models of the ventral visual stream: (i) they can predict image recognition difficulty well for humans (Fig. 1A), and (ii) they can be used to perturb images in a way that enhances human ability to accurately report the original ground truth (Fig. 1B). In other words, these models can be used out-of-the-box as category-recognition difficulty estimators, and category-percept enhancers. As such, we propose using them to augment a visual learning sequence that is used to learn and practice recognition of unfamiliar image classes in various challenging categorization tasks.
Fig. 2 summarizes our approach. In the naive baseline scenario, the novice learner is presented with a sequence of training trials, where in each, a randomly selected image from one of $N$ categories is presented, and the user is tasked to perform an $N$-way categorization task through an online survey platform. During this training phase, the learner receives correct/incorrect feedback after each trial. Notably, the category labels are made meaningless by assigning them to be a random Greek name unrelated to the task, such that learners start at chance level. At the end of the training phase, the experiment transitions into the test phase, where the same task continues over a held-out set of images, and no feedback is delivered. During the test phase we measure the visual learning outcome of interest, the test accuracy (Fig. 2A-C).
Harnessing the key observations on robustified ANNs, our approach uses a model to optimize the training phase in a way that improves the test accuracy via two mechanisms: (i) sampling training images based on their predicted recognition difficulty, and (ii) enhancing the training images. Both mechanisms have “strength” control knobs which are generally time dependent, namely the maximally allowable difficulty of an image at a given trial, and the corresponding enhancement strength. The latter is approximated via the $\ell_{2}$-norm pixel-budget $\epsilon$. Using these mechanisms, the user of our approach can flexibly define arbitrary time-dependent profiles for image selection and enhancement (Fig. 2D-F). In this study, we focused on a linear ramp profile for the allowable image difficulty at a given time and on exponential tapering of the enhancement $\epsilon$. Intuitively, this should correspond to an easy-to-challenging traversal during the training phase. Notably, extensively optimizing these profiles was not our goal.
Incorporating these mechanisms led to significant gains in the test-time accuracy of human participants, while also requiring less time to complete the training phase (which includes a constant number of trials). This result was robustly obtainable across the varied image domains and category spaces tested. In particular, we considered three categorization tasks: moth species in natural images, dermoscopy images (skin lesion diagnosis), and histology images (benign lesions vs pre-cancerouss lesion diagnosis). We next describe the model and explain the two model-based mechanisms that drive the boost in visual learning.

### v2

# L-WISE: Boosting human visual category learning through model-based image selection and enhancement

@@ Abstract
The currently leading artificial neural network models of the visual ventral stream – which are derived from a combination of performance optimization and robustification methods – have demonstrated a remarkable degree of behavioral alignment with humans on visual categorization tasks. We show that image perturbations generated by these models can enhance the ability of humans to accurately report the ground truth class. Furthermore, we find that the same models can also be used out-of-the-box to predict the proportion of correct human responses to individual images, providing a simple, human-aligned estimator of the relative difficulty of each image. Motivated by these observations, we propose to augment visual learning in humans in a way that improves human categorization accuracy at test time. Our learning augmentation approach consists of (i) selecting images based on their model-estimated recognition difficulty, and (ii) applying image perturbations that aid recognition for novice learners. We find that combining these model-based strategies leads to categorization accuracy gains of 33-72% relative to control subjects without these interventions, on unmodified, randomly selected held-out test images. Beyond the accuracy gain, the training time for the augmented learning group was also shortened by 20-23%, despite both groups completing the same number of training trials. We demonstrate the efficacy of our approach in a fine-grained categorization task with natural images, as well as two tasks in clinically relevant image domains – histology and dermoscopy – where visual learning is notoriously challenging. To the best of our knowledge, our work is the first application of artificial neural networks to increase visual learning performance in humans by enhancing category-specific image features.
@@ 1 Introduction
Figure 1: Robustified ANNs can be used out-of-the-box as image recognition difficulty estimators and ground truth percept enhancers. We consider a 16-way basic animal classification task. Panel A1 shows the correspondence between human categorization accuracy and model-computed ground truth logit activation values. The curve denotes a logistic regression model predicting the probability of a correct response using only the logit value ($p<0.001$ from the Wald statistic, $\text{AUC}=0.72$ under 10-fold cross validation). A2 shows example images with varying ground truth logit values (predicted difficulty). B1 shows how perturbing images via ground truth logit maximization increases human recognition accuracy progressively with the $\ell_{2}$-norm perturbation pixel budget $\epsilon$. Other off-the-shelf image enhancement methods do not increase categorization accuracy, despite inducing larger perturbations of $\epsilon=43$, $\epsilon=106$, and $\epsilon=26$ on average from left to right. B2 shows example images: unmodified (left), enhanced by ground truth logit maximization with pixel budgets ${\epsilon}=10$ and ${\epsilon}=20$ (middle), and enhanced by baseline off-the-shelf methods (right). All vertical error bars are 95% confidence intervals by bootstrap. Horizontal error bars (panel A1) show the standard deviation among images within each logit value bin.
@@
Figure 2: Robustified ANNs can be used to boost image category learning in humans. A novice human learner undertakes a challenging image categorization task, which consists of a training phase (B) and a test phase (C). Images for both phases are randomly drawn from a labeled image dataset of unfamiliar fine-grained categories (A). Feedback (correct/incorrect, with indication of the correct category) is delivered after each trial during the training phase only. Our proposed “Logit-Weighted Image Selection and Enhancement” (L-WISE) approach uses an ANN model (D) to augment the visual curriculum by using the difficulty score to sample images based on a predefined increasing schedule of maximal difficulty per trial (E), and by enhancing images for easier recognition with an enhancement magnitude that decreases along a predefined schedule (F).
@@
Over the last decade, artificial neural network (ANN) models have demonstrated superior performance as image-computable emulators of neural processing along the human and monkey ventral visual stream. Iterative efforts have developed models that are increasingly aligned with primate vision, as measured by their ability to predict both neural activity and behavioral responses (Schrimpf et al., 2018). Beyond prediction, a promising class of these models – “robustified” deep ANNs  (Ma̧dry et al., 2018) – has been shown to enable the generation of image perturbations that predictably control both ventral stream neural activity (Guo et al., 2022) and human object categorization reports (Gaziv et al., 2023; Croce & Hein, 2020). In our work, we ask whether the prediction and control capabilities of robustified ANNs can be used to enhance human performance at visual categorization tasks.
Beyond categorization of familiar visual categories, one practically important task is learning to recognize new, unfamiliar categories. Although humans can readily learn many new categories even from a single example (Lake et al., 2011), some consequential tasks require extensive training to reach high levels of performance. For example, medical specialists such as pathologists and radiologists devote numerous hours to mastering the diagnosis of various diseases from medical images. We identify a potential strategy to accelerate the visual learning process by extrapolating from the perceptual learning literature. Simple visual tasks, such as line orientation discrimination, reveal a curriculum effect whereby providing easy examples to a novice human learner before gradually increasing the difficulty promotes faster perceptual learning (Lu & Dosher, 2022). Motivated by these findings, we ask whether the human-aligned nature of robustified ANNs allows them to augment human learning in complex image domains, chiefly by reducing the initial task difficulty and then increasing the difficulty as learning progresses. A demonstration that these models can be used to enhance learning serves as an additional scientific test of the models’ alignment with human perception, while also pioneering a potentially beneficial application of ANNs in education.
@@
To test the viability of enhancing image category learning in humans, we first establish two key empirical observations, summarized in Fig. 1: (i) we found that the human error rate in an image categorization task is strongly predicted by the ground truth logit activation of a robustified ANN, making it a valid image recognition difficulty score for humans; (ii) we found that this relationship also holds in reverse – pixel-level perturbations can be generated by the model to maximize the ground truth logit activation, producing an “enhanced” version of the image that is easier for humans to recognize as the ground truth category. While previous findings demonstrate that model-guided perturbations can modulate human perception away from the ground truth category (Gaziv et al., 2023), we conversely seek to amplify category-relevant features to facilitate correct classification. We observe that our model-based image enhancements yield significant increases in human accuracy on a classification task with basic animal categories, unlike conventional image enhancement techniques that adjust low-level image properties such as lighting and color-contrast. We thus propose an algorithm that combines model-based image enhancement and image difficulty prediction to generate optimized curricula for novice humans learning challenging image categorization tasks.
@@
Our proposed method, “Logit-Weighted Image Selection and Enhancement” (L-WISE), is illustrated in Fig. 2. In our primary experimental setting, a human participant learns an unfamiliar image categorization task by viewing a series of examples, providing a category judgment on each image before receiving feedback indicating the correct category (Fig. 2A,B). Upon completion of the training phase, test accuracy is measured on held-out images in similar trials without feedback (Fig. 2C). L-WISE intervenes on this naive visual learning baseline by using a robustified ANN model in two ways: (i) it uses the ground truth logit difficulty score to sample training images based on a predefined, increasing schedule of maximal difficulty per trial (Fig. 2D,E); (ii) it enhances training images with a perturbation magnitude that decreases along a similarly predefined schedule (Fig. 2D,F).
@@
Despite the human visual system being well-adapted for rapidly learning new visual categories, we find that L-WISE gives rise to substantial accuracy gains of 33-72% in test-time accuracy margins above chance relative to control participants. In addition to improved accuracy, L-WISE also significantly reduces the training phase duration by 20-23% with a constant number of trials. We demonstrate these effects across three varied image domains and category spaces: moth species in natural photographs, skin lesions in dermoscopy images, and pathologic findings in histology images.
Our main contributions are:
@@
 $\bullet$ We establish a new state-of-the-art in predicting image recognition difficulty for humans, using the robustified ANN ground truth logit activation value as a simple but effective difficulty metric.
@@
 $\bullet$ We show that robustified ANNs can guide image perturbations that enhance the ability of humans to accurately report the associated ground truth category label.
@@
 $\bullet$ We propose a novel model-based visual learning augmentation approach for humans that substantially increases test-time categorization accuracy and also reduces training time. To the best of our knowledge, this is the first application of image enhancement to augment human visual learning.
@@
 $\bullet$ We demonstrate the broad applicability of our proposed method in a variety of image domains, including clinically relevant dermoscopy and histology categorization tasks.
@@ 2 Related Work
We develop two important capabilities that form the foundation of our approach to assisting human learners: (1) state-of-the-art predictions of the recognition difficulty of images for humans, and (2) image perturbations that increase human categorization accuracy. Many works have ranked the difficulty of images to design curricula for training ANNs (Wang et al., 2021b). Leading approaches include the c-score learning speed proxy (Jiang et al., 2021) and the prediction depth (Baldock et al., 2021) calculated for each image. Mayo et al. (2023) applied both of these techniques to predict the recognition difficulty of natural images for humans, defined as either the minimum viewing time required to classify a given image correctly, or (as in our work) the proportion of humans who correctly classify it. Here, we show that a robustified ANN model’s logit score associated with the ground truth class is a more accurate predictor of image difficulty for humans than prior methods.
@@
Enhancing image quality has been the focus of many previous studies (Qi et al., 2021), ranging from correction of low-level properties such as lighting and contrast (e.g., Zuiderveld (1994); Jobson et al. (1997)) to ANN models that “upsample” images to higher resolutions (Anwar et al., 2020). However, very little research has focused on enhancing images to more strongly represent a specific category. Previous works in this vein focused on making images easier for ANN models to correctly classify (Kim et al., 2023; Tussupov et al., 2023) or less vulnerable to subsequent adversarial attacks (Salman et al., 2021; Frosio & Kautz, 2023). However, such methods do not strongly affect human responses due to perceptual misalignment between humans and naive ANNs (Gaziv et al., 2023).
@@
Many studies have focused on model-human alignment. Brain-Score benchmarks models in terms of neural representations and behavior (Schrimpf et al., 2018). “Harmonization” methods drive alignment using an auxiliary objective on ANN-predicted feature importance maps and crowd-sourced human maps (Fel et al., 2022). Other works introduce architecture components to account for additional aspects of human vision, such as the dorsal-stream pathway (Choi et al., 2023), or recurrent connections for contextual reasoning (Bomatter et al., 2021) and visual search (Gupta et al., 2021).
@@
A key property that enables ANNs to generate human-interpretable image perturbations is that of perceptually aligned gradients, which is closely related to adversarial robustness and can be induced through adversarial training (Ganz et al., 2023; Gaziv et al., 2023). In our present work, we apply adversarially-trained ANNs to enhance images such that they are more strongly associated with their ground truth label by the guiding model and by humans. To the best of our knowledge, we are the first to demonstrate improved human performance on image classification tasks through category-specific image enhancement.
@@
Our primary goal is to apply difficulty prediction and image enhancement to augment human learning. The emerging field of machine teaching (Zhu, 2015) employs machine learning to find or generate optimal “teaching sets” that can be used to train models or humans. While many such approaches have been successfully applied to training machine learning models (e.g., Liu et al. (2017); Qiu et al. (2023)), few studies have successfully enhanced image category learning in humans and most of these focus on teaching set selection. Singla et al. (2014) propose STRICT, which optimizes the expected decrease in learner error based on how the selected images and their labels constrain a linear hypothesis class in a feature space. Johns et al. (2015) extend a similar approach to select images in an online fashion by modeling the learner’s progress. MaxGrad (Wang et al., 2021a) uses bi-level optimization to iteratively refine a teaching set by modeling learners as optimal empirical risk minimizers. Most similar to our work are approaches like EXPLAIN (Mac Aodha et al., 2018), which uses ANN class activation maps (CAMs) to highlight relevant image regions while providing feedback to the learner. EXPLAIN also selects a curriculum of images based on (a) a multi-class adaptation of STRICT, (b) representativeness (mean feature-space distance to other images of the same class), and (c) the estimated difficulty (entropy) of the CAM explanations. Chang et al. (2023) use bounding boxes to highlight image regions attended to by experts and not novices, allowing humans to more accurately match bird or flower images to one species among five shown in a gallery.
@@
Our approach to learning augmentation departs from previous studies in several ways. We make explicit estimates of image difficulty with unprecedented accuracy to select easier images for early-stage learners. Our work is unique in employing category-specific image enhancement, which is a novel technique in itself, to improve the teaching efficacy of a given set of images. While Mac Aodha et al. (2018) and Chang et al. (2023) help learners by explicitly highlighting where learners should attend to in each image, we take a distinct and complementary approach by implicitly highlighting what learners must attend to in order to classify images correctly.
@@ 3 Overview of approach and experiments
Our approach to improving visual learning in humans is based on two key observations regarding human-aligned ANN models of the ventral visual stream: (i) they can accurately predict the recognition difficulty of specific images for humans (Fig. 1A), and (ii) they can be used to perturb images in a way that enhances the ability of humans to accurately report the original ground truth category (Fig. 1B). In other words, these models can be used out-of-the-box as category-recognition difficulty estimators, and category-percept enhancers. As such, we propose using them to design sequences of images that humans can use to more efficiently learn to recognize unfamiliar image categories. We test our approach in a variety of challenging image domains: natural photographs (moth species classification), dermoscopy images (skin lesion classification), and histology images (benign vs. pre-cancerous classification).